• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 02:12:01 on December 20,1998

    Entry number 8941

    Followups:

    keyword=NMR /rotation to perpendicular

    01:34 NMR 39.42 % (H20)
    01:36 rotate to perpendicular


    Perpendicular laser #6 will not stay on. The temperature seems to be very high
    (30 C as compared to 15 C on all the others). We are forced to run with
    2 perpendicular lasers.

    --------------------------------------
    01:28 NMR 3// lasers
    01:29 lasers off
    01:32 target to 50545
    01:36 rotation complete
    01:36 laser #6 on then off
    01:39 laser 7 on
    01:40 laser 6 on then off
    01:41 laser 8 on